OpenStudy (dragons1415):

So... Any Alg. 2 people willing to help? :P

OpenStudy (confluxepic):

Sure.

OpenStudy (confluxepic):

I can try.

OpenStudy (dragons1415):

Given that f(x) = 2x + 5, find -2* f(x + 1). Here's the given hint: Find f(x + 1) than multiply by -2.

OpenStudy (skyz):

Wrong account o-o

OpenStudy (skyz):

Yeah this is to strong for what i have

OpenStudy (dragons1415):

Lol Skyz xD

OpenStudy (anonymous):

dragons, do you know what f(x+1) means?

OpenStudy (skyz):

Do you even remember me~

OpenStudy (dragons1415):

function of x+1 bagajr .-.

OpenStudy (dragons1415):

Yeah, of course I do, Skyz!

OpenStudy (confluxepic):

f(x) is y.

OpenStudy (anonymous):

So plug in x+1 into f(x) to get f(x+1)

OpenStudy (anonymous):

Do you know what I mean?

OpenStudy (confluxepic):

f(x+1)=2(x+1)+5

OpenStudy (dragons1415):

...... Conflux, I finally realized what I've been doing wrong ._.

OpenStudy (confluxepic):

(2(x+1)+5)=2x+2+5

OpenStudy (confluxepic):

2x+2+5=2x+7

OpenStudy (confluxepic):

(2x+7)(-2)=-4x-14

OpenStudy (dragons1415):

Thanks guys! <3

OpenStudy (confluxepic):

Thanks for the medal.

OpenStudy (dragons1415):

No prob .-.

OpenStudy (dragons1415):

Wish I could give out more medals though...